LO 2.3Markson and Sons leases a copy machine with terms that include a fixed fee each month of $500 plus a charge for each copy made. The company uses the high-low method to analyze costs. If Markson paid $360 for 5.000 copies and $280 for 3,000 copies how much would Markson pay if it made 7.500 copies ?
Answer #1

Variable cost per copy = Change in Cost/Change in Copy = (360-280)/(5000-3000) = 0.04

Fixed cost = 360-(5000*.04) = 160

Markson should pay = 500+(7500*.04) = 800